# Space-time FEM-BEM couplings for parabolic transmission problems

### v1

@@ 1.1. Model problem
For $d\geq 2$, let $\Omega\subset\mathbb{R}^{d}$ be a bounded Lipschitz domain with connected boundary $\Gamma$ and $T>0$ a given final time with corresponding time interval $I:=(0,T)$. We abbreviate the space-time cylinder by $Q:=I\times\Omega$, and denote its lateral boundary by $\Sigma:=I\times\Gamma$. The corresponding outward normal vector on $\Sigma$ is denoted by ${\boldsymbol{n}}=(0,{\boldsymbol{n}}_{\boldsymbol{x}})\in\mathbb{R}^{d+1}$. Moreover, let $\Omega^{\rm ext}:=\mathbb{R}^{d}\setminus\overline{\Omega}$ and $Q^{\rm ext}:=I\times\Omega^{\rm ext}$ be the corresponding exterior domains. For ${\boldsymbol{A}}\in L^{\infty}(Q)^{d\times d}$ symmetric uniformly positive definite and given data $f:Q\to\mathbb{R}$, $u_{0}:\Omega\to\mathbb{R}$, $u_{D}:\Sigma\to\mathbb{R}$, and $\phi_{N}:\Sigma\to\mathbb{R}$, whose properties will be specified below, we consider the parabolic transmission problem to find $u:Q\rightarrow\mathbb{R}$ and $u^{\rm ext}:Q^{\rm ext}\rightarrow\mathbb{R}$ such that
@@ 1.2. Motivation and existing works
Problem (1.1) models, for instance, heat conduction in a heterogeneous body $\Omega$ that radiates heat into the surrounding homogeneous medium $\Omega^{\rm ext}$. In this context, the unbounded domain $\Omega^{\rm ext}$ can of course be viewed as a simplification of a surrounding bounded $\widetilde{\Omega}$ that is substantially larger than $\Omega$. The work [42] provides a concrete interpretation of (1.1) in terms of photothermal spectroscopy. Other works that investigate the heat equation on unbounded domains include for instance [4, 44, 57]. Assuming that ${\boldsymbol{A}}$ is scalar and constant on $Q$, $f=0$ on $Q$, and $u_{0}=0$ on $\Omega$, [41] discretize (1.1) via a boundary element method (BEM) in the interior and the exterior, and the resulting semidiscrete system is solved by convolution quadrature [31, 33, 32] in time. In the PhD thesis [8], a coupling of the space-time finite element method (FEM) [50] based on some Hilbert-type transformation and space-time BEM [1, 39, 5] is proposed, and a numerical example for $d=1$ is presented. However, no numerical analysis of the scheme is given, and it is unclear whether the resulting discrete systems are inf-sup stable or even uniquely solvable. In addition, the computation of the mentioned Hilbert-type transformation is possible but non-trivial for higher dimensions [51, 54, 55]. A provably stable numerical scheme that approximately solves (1.1) in its general form has been missing in the literature up to now. It is the objective of this manuscript to introduce and analyze such a method.
@@ 1.4. Space-time discretization
While a time-stepping method as [41] only needs to solve relatively small linear systems in each time step, a potential downside of simultaneous space-time methods [1, 39, 5, 46, 49, 9, 50, 20, 24, 25, 17] is that one has to solve one huge linear system at once. In addition, boundary element discretizations lead to dense matrices due to the nonlocality of the boundary integral operators. On the other hand, simultaneous space-time methods in principle have the potential for massive parallelization, see, e.g., [23, 38, 52] for FEM and [10, 56] for BEM. Preconditioners for the employed space-time FOSLS formulation [20, 24] that can be applied in parallel are still missing in the literature and shall be addressed in the future. Having efficient preconditioners for FEM and BEM at hand, block-diagional preconditioners give efficient preconditioners for FEM-BEM couplings as has been analyzed in, e.g., [37, 15] for elliptic transmission problems. Moreover, the dense BEM matrices can be compressed via the fast multipole method and $\mathcal{H}$-matrices, cf. [35, 36, 29, 53].
@@ 2.1. Sobolev spaces
from, e.g., [7, pages 480 & 494]. Next, we recall that the restriction $(\cdot)|_{\Sigma}$ is a bounded linear mapping from $H^{1,1/2}(Q)$ to $H^{1/2,1/4}(\Sigma)$; see, e.g., [30, Chapter 4, Theorem 2.1]. Finally, [5, Proposition 2.18] states that the normal derivative $\nabla_{\boldsymbol{x}}(\cdot)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$ is a bounded linear mapping from the space $\big{\{}u\in H^{1,1/2}(Q)\,:\,\partial_{t}u-\Delta_{\boldsymbol{x}}u\in L^{2}(Q)\big{\}}$ equipped with the canonical graph norm to $H^{-1/2,-1/4}(\Sigma)$.
For $R>0$, let $B_{R}(\boldsymbol{0}):=\big{\{}{\boldsymbol{x}}\in\mathbb{R}^{d}\,:\,|{\boldsymbol{x}}|<R\big{\}}$ with boundary $\Gamma_{R}:=\partial B_{R}(\boldsymbol{0})$. For $R>0$ such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$, we abbreviate $\Omega^{\rm ext}_{R}:=\Omega^{\rm ext}\cap B_{R}(\boldsymbol{0})$, $Q_{R}^{\rm ext}:=I\times\Omega_{R}^{\rm ext}$, and $\Sigma_{R}:=I\times\Gamma_{R}$ with corresponding outward normal vector ${\boldsymbol{n}}_{R}=(0,{\boldsymbol{n}}_{R,{\boldsymbol{x}}})$. We define the Sobolev spaces on $Q_{R}^{\rm ext}$ and $\Sigma_{R}$ as for $Q$ and $\Sigma$.
@@ 2.2. Boundary integral operators
These potentials can be extended to bounded linear mappings $\widetilde{\mathcal{V}}:H^{-1/2,-1/4}(\Sigma)\to H^{1,1/2}(Q)$ and $\widetilde{\mathcal{K}}:H^{1/2,1/4}(\Sigma)\to H^{1,1/2}(Q)$. For any $\phi\in H^{-1/2,-1/4}(\Sigma)$ and $u\in H^{1/2,1/4}(\Sigma)$, $\widetilde{\mathcal{V}}\phi$ and $\widetilde{\mathcal{K}}u$ are indeed potentials in the sense that they satisfy the heat equation. This also readily implies that $\widetilde{\mathcal{V}}\phi,\widetilde{\mathcal{K}}u\in L^{2}(I;H^{1}(\Omega)\cap H^{1}(I;H^{-1}(\Omega))$. Finally, the restriction of $\widetilde{\mathcal{V}}\phi$ and $\widetilde{\mathcal{K}}u$ to $\{0\}\times\Omega$ is zero, i.e., $(\widetilde{\mathcal{V}}\phi)(0,\cdot)=(\widetilde{\mathcal{K}}u)(0,\cdot)=0$. Conversely, if $u\in L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ is a weak solution of the heat equation $\partial_{t}u-\Delta_{\boldsymbol{x}}u=0$ with $u(0,\cdot)=0$, then there holds the representation formula
@@
We further define the single-layer operator $\mathcal{V}$ and the double-layer operator $\mathcal{K}$ as in (2.2)–(2.3) with $(t,{\boldsymbol{x}})\in I\times\Sigma$ instead of $(t,{\boldsymbol{x}})\in Q$. These operators on the lateral boundary $\Sigma$ can be extended to bounded linear mappings $\mathcal{V}:H^{-1/2,-1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$ and $\mathcal{K}:H^{1/2,1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$. In addition, $\mathcal{V}$ is coercive, i.e.,
@@
$({\mathcal{W}}u)(t,{\boldsymbol{x}})$ $:=-\int_{\Sigma}\nabla_{\boldsymbol{x}}(\nabla_{\boldsymbol{y}}G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})\cdot{\boldsymbol{n}}_{\boldsymbol{y}})\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\,u(s,{\boldsymbol{y}})\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\quad\text{for all }(t,{\boldsymbol{x}})\in\Sigma.$ (2.9)
These operators can be extended to bounded linear mappings $\mathcal{W}:H^{1/2,1/4}(\Sigma)\to H^{-1/2,-1/4}(\Sigma)$ and $\mathcal{N}:H^{-1/2,-1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$. In addition, $\mathcal{W}$ is coercive, i.e.,
@@
As a matter of fact, [5, Theorem 3.11] even states the following coercivity
@@ 2.3. Weak formulation
Given right-hand sides $f\in L^{2}(Q)$, $u_{0}\in L^{2}(\Omega)$, $u_{D}\in H^{1/2,1/4}(\Sigma)$, and $\phi_{N}\in H^{-1/2,-1/4}(\Sigma)$, we look for weak solutions of (1.1), i.e., $u\in L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ and $u^{\rm ext}|_{Q^{\rm ext}_{R}}\in L^{2}(I;H^{1}(\Omega^{\rm ext}_{R}))$ for all $R>0$ such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$. Note that $\partial_{t}u^{\rm ext}-\Delta_{\boldsymbol{x}}u^{\rm ext}=0$ readily yields the additional regularity $u^{\rm ext}|_{Q^{\rm ext}_{R}}\in L^{2}(I;H^{1}(\Omega^{\rm ext}_{R}))\cap H^{1}(I;H^{-1}(\Omega^{\rm ext}_{R}))$; see [5, Lemma 2.15]. By the properties of the involved Sobolev spaces mentioned in Section 2.1, the left-hand sides in (1.1) are indeed all well defined.
Let $R>0$ be sufficiently large such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$. With the associated single- and double-layer potential $\widetilde{\mathcal{V}}_{R}$ and $\widetilde{\mathcal{K}}_{R}$ on $\Sigma_{R}$, the representation formula (2.4) applied to $Q^{\rm ext}_{R}$ instead of $Q$ yields that any weak solution of the exterior problem (1.1b)&(1.1d) is given by
@@
In what follows, we show that the radiation condition (1.1g) implies that the last two terms in (2.14) vanish as $R\to\infty$. We start with the following auxiliary result, which provides a crude estimate for the gradient $\nabla_{\boldsymbol{x}}u^{\rm ext}$.
@@
By parabolic regularity, e.g., [16, Theorem 1.58], the homogeneous heat equation (1.1b) implies that $u^{\rm ext}$ is even a smooth strong solution of $\partial_{t}u^{\rm ext}-\Delta_{\boldsymbol{x}}u^{\rm ext}=0$ on $Q^{\rm ext}$. In particular, each component of $\nabla_{\boldsymbol{x}}u^{\rm ext}$ satisfies the homogeneous heat equation strongly on $Q^{\rm ext}$. We may thus apply the mean-value property for the heat equation, e.g., [14, Section 2.3, Theorem 3] for each component: Let $R>0$ be sufficiently large such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$ and $|u^{\rm ext}(t,{\boldsymbol{x}})|\leq 1$ for all $t\in I$ and all ${\boldsymbol{x}}\in\Omega^{\rm ext}$ with $|{\boldsymbol{x}}|\geq R$. Now, we fix coordinates $t\in I$ and ${\boldsymbol{x}}\in\Omega^{\rm ext}$ with $|{\boldsymbol{x}}|\geq 2R$. For $r>0$, we define the corresponding heat ball
@@
For sufficiently small $r>0$ with $E(t,{\boldsymbol{x}};r)\subset Q^{\rm ext}$, it holds the mean-value property
@@
If $d=3$, the last integral is just $\tfrac{1}{2\pi}t$, which concludes the proof in this case even for $\epsilon=0$. If $d=2$, we write $\ln\frac{t}{4\pi s}=\ln t-\ln 4\pi-\ln s\leq\ln T-\ln s\leq 2\max\{|\ln T|,|\ln s|\}$ and get that
$|\nabla_{\boldsymbol{x}}u^{\rm ext}(t,{\boldsymbol{x}})|\lesssim\frac{1}{t}\int_{0}^{\frac{1}{4\pi}t}s^{-1/2}\Big{(}|\ln T|^{3/2}+|\ln s|^{3/2}\Big{)}\,{\rm d}s\lesssim t^{-1/2-\epsilon}.$
@@
It remains to show that the last two terms in (2.14) tend to zero. To this end, let $(t,{\boldsymbol{x}})\in Q_{R}^{\rm ext}$ be arbitrary but fixed. The simple estimate $|G(s,{\boldsymbol{y}})|\lesssim s^{1/2}|{\boldsymbol{y}}|^{-d-1}$ for all $s\geq 0$, ${\boldsymbol{y}}\in\mathbb{R}^{d}$ (see, e.g., the proof of [5, Theorem 3.11]), the preceding lemma, and $|\Gamma_{R}|\simeq R^{d-1}$ show for sufficiently large $R>0$ and $20<\epsilon<1/$ 2 that
@@
Conversely, in the proof of [5, Theorem 3.11], it is shown that any function of the form
@@
To be more precise, the cited [5, Theorem 3.11] states for $u^{\rm ext}$ defined as in (2.20) that $u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-d})$ and $\nabla_{\boldsymbol{x}}u^{\rm ext}(\cdot,{\boldsymbol{x}})\cdot{\boldsymbol{x}}={\mathcal{O}}(|x|^{-d})$ as $|x|\to\infty$. Using the same argument, one even sees that $u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-\mu})$ and $\nabla_{\boldsymbol{x}}u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-\mu})$ as $|x|\to\infty$ for all fixed $\mu\geq 0$ (with hidden constants depending only on $d$, $\mu$, $u^{\rm ext}$, and $T$). The assumption (1.1g) that $u^{\rm ext}$ converges to zero as $|{\boldsymbol{x}}|\to\infty$ thus automatically yields a much stronger decay of $u^{\rm ext}$ and its spatial gradient $\nabla_{\boldsymbol{x}}u^{\rm ext}$. On the other hand, to guarantee that the last two terms in (2.14) converge to zero, it is clear from our proof that one could even allow for a certain growth of $u^{\rm ext}$ (and $\nabla_{\boldsymbol{x}}u^{\rm ext}$). Here and in the following, the minimal required radiation condition would be to explicitly assume that those two terms vanish as $R\to\infty$.

### v2

@@ 1.1. Model problem
For $d\in\{2,3\}$, let $\Omega\subset\mathbb{R}^{d}$ be a bounded Lipschitz domain with connected boundary $\Gamma$ and $T>0$ a given final time with corresponding time interval $I:=(0,T)$. We abbreviate the space-time cylinder by $Q:=I\times\Omega$, and denote its lateral boundary by $\Sigma:=I\times\Gamma$. The corresponding outward normal vector on $\Sigma$ is denoted by ${\boldsymbol{n}}^{\top}=(0,{\boldsymbol{n}}_{\boldsymbol{x}}^{\top})\in\mathbb{R}^{d+1}$. Moreover, let $\Omega^{\rm ext}:=\mathbb{R}^{d}\setminus\overline{\Omega}$ and $Q^{\rm ext}:=I\times\Omega^{\rm ext}$ be the corresponding exterior domains. For ${\boldsymbol{A}}\in L^{\infty}(Q)^{d\times d}$ symmetric uniformly positive definite and given data $f:Q\to\mathbb{R}$, $u_{0}:\Omega\to\mathbb{R}$, $u_{D}:\Sigma\to\mathbb{R}$, and $\phi_{N}:\Sigma\to\mathbb{R}$, whose properties will be specified below, we consider the parabolic transmission problem to find $u:Q\rightarrow\mathbb{R}$ and $u^{\rm ext}:Q^{\rm ext}\rightarrow\mathbb{R}$ such that
@@ 1.2. Motivation and existing works
Problem (1.1) models, for instance, heat conduction in a heterogeneous body $\Omega$ that radiates heat into the surrounding homogeneous medium $\Omega^{\rm ext}$. In this context, the unbounded domain $\Omega^{\rm ext}$ can of course be viewed as a simplification of a surrounding bounded $\widetilde{\Omega}$ that is substantially larger than $\Omega$, and we refer for instance to [6, 36, 20, 33, 4] for parabolic transmission problems on bounded domains. The work [50] provides a concrete interpretation of (1.1) in terms of photothermal spectroscopy. Other works that investigate the heat equation on unbounded domains include for instance [13, 5, 34, 37, 59, 52, 63]. Assuming that ${\boldsymbol{A}}$ is scalar and constant on $Q$, $f=0$ on $Q$, and $u_{0}=0$ on $\Omega$, [49] discretize (1.1) via a boundary element method (BEM) in the interior and the exterior, and the resulting semidiscrete system is solved by convolution quadrature [39, 41, 40] in time. In the PhD thesis [10], a coupling of the space-time finite element method (FEM) [58] based on some Hilbert-type transformation and space-time BEM [1, 47, 7] is proposed, and a numerical example for $d=1$ is presented. However, no numerical analysis of the scheme is given, and it is unclear whether the resulting discrete systems are inf-sup stable or even uniquely solvable. A provably stable numerical scheme that approximately solves (1.1) in its general form has been missing in the literature up to now. It is the objective of this manuscript to introduce and analyze such a method.
@@ 1.4. Space-time discretization
While a time-stepping method as [49] only needs to solve relatively small linear systems in each time step, a potential downside of simultaneous space-time methods [1, 47, 7, 54, 57, 11, 58, 24, 28, 29, 21] is that one has to solve one huge linear system at once. In addition, boundary element discretizations lead to dense matrices due to the nonlocality of the boundary integral operators. On the other hand, simultaneous space-time methods in principle have the potential for massive parallelization, see, e.g., [27, 46, 60] for FEM and [12, 62] for BEM. Preconditioners for the employed space-time FOSLS formulation [24, 28] that can be applied in parallel are still missing in the literature and shall be addressed in the future. Having efficient preconditioners for FEM and BEM at hand, block-diagonal preconditioners give efficient preconditioners for FEM-BEM couplings as has been analyzed in, e.g., [45, 18] for elliptic transmission problems. Moreover, the dense BEM matrices can be compressed via the fast multipole method and $\mathcal{H}$-matrices, cf. [43, 44, 35, 61].
@@ 2.1. Sobolev spaces
from, e.g., [9, pages 480 & 494]. Next, we recall that the restriction $(\cdot)|_{\Sigma}$ is a bounded linear mapping from $H^{1,1/2}(Q)$ to $H^{1/2,1/4}(\Sigma)$; see, e.g., [38, Chapter 4, Theorem 2.1]. Finally, [7, Proposition 2.18] states that the normal derivative $\nabla_{\boldsymbol{x}}(\cdot)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$ is a bounded linear mapping from the space $\big{\{}u\in H^{1,1/2}(Q)\,:\,\partial_{t}u-\Delta_{\boldsymbol{x}}u\in L^{2}(Q)\big{\}}$ equipped with the canonical graph norm to $H^{-1/2,-1/4}(\Sigma)$.
For $R>0$, let $B_{R}(\boldsymbol{0}):=\big{\{}{\boldsymbol{x}}\in\mathbb{R}^{d}\,:\,|{\boldsymbol{x}}|<R\big{\}}$ with boundary $\Gamma_{R}:=\partial B_{R}(\boldsymbol{0})$. For $R>0$ such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$, we abbreviate $\Omega^{\rm ext}_{R}:=\Omega^{\rm ext}\cap B_{R}(\boldsymbol{0})$, $Q_{R}^{\rm ext}:=I\times\Omega_{R}^{\rm ext}$, and $\Sigma_{R}:=I\times\Gamma_{R}$ with corresponding outward normal vector ${\boldsymbol{n}}_{R}^{\top}=(0,{\boldsymbol{n}}_{R,{\boldsymbol{x}}}^{\top})$. We define the Sobolev spaces on $Q_{R}^{\rm ext}$ and $\Sigma_{R}$ as for $Q$ and $\Sigma$.
@@ 2.2. Boundary integral operators
These potentials can be extended to bounded linear mappings $\widetilde{\mathcal{V}}:H^{-1/2,-1/4}(\Sigma)\to H^{1,1/2}(Q)$ and $\widetilde{\mathcal{K}}:H^{1/2,1/4}(\Sigma)\to H^{1,1/2}(Q)$. For any $\phi\in H^{-1/2,-1/4}(\Sigma)$ and $u\in H^{1/2,1/4}(\Sigma)$, $\widetilde{\mathcal{V}}\phi$ and $\widetilde{\mathcal{K}}u$ are indeed potentials in the sense that they satisfy the homogeneous heat equation. This also readily implies that $\widetilde{\mathcal{V}}\phi,\widetilde{\mathcal{K}}u\in L^{2}(I;H^{1}(\Omega)\cap H^{1}(I;H^{-1}(\Omega))$. Finally, the restriction of $\widetilde{\mathcal{V}}\phi$ and $\widetilde{\mathcal{K}}u$ to $\{0\}\times\Omega$ is zero, i.e., $(\widetilde{\mathcal{V}}\phi)(0,\cdot)=(\widetilde{\mathcal{K}}u)(0,\cdot)=0$. Conversely, if $u\in L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ is a weak solution of the heat equation $\partial_{t}u-\Delta_{\boldsymbol{x}}u=0$ with $u(0,\cdot)=0$, then there holds the representation formula
@@
We further define the single-layer operator $\mathcal{V}$ and the double-layer operator $\mathcal{K}$ as in (2.2)–(2.3) with $(t,{\boldsymbol{x}})\in I\times\Sigma$ instead of $(t,{\boldsymbol{x}})\in Q$. Note that the corresponding integral are now (weakly) singular. These operators on the lateral boundary $\Sigma$ can be extended to bounded linear mappings $\mathcal{V}:H^{-1/2,-1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$ and $\mathcal{K}:H^{1/2,1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$. In addition, $\mathcal{V}$ is coercive, i.e.,
@@
$({\mathcal{W}}u)(t,{\boldsymbol{x}})$ $:=-{\boldsymbol{n}}_{\boldsymbol{x}}\cdot\nabla_{\boldsymbol{x}}\int_{\Sigma}\nabla_{\boldsymbol{y}}G(t-s,{\boldsymbol{x}}-{\boldsymbol{y}})\cdot{\boldsymbol{n}}_{\boldsymbol{y}})u(s,{\boldsymbol{y}})\,{\rm d}{\boldsymbol{y}}\,{\rm d}s\quad\text{for all }(t,{\boldsymbol{x}})\in\Sigma.$ (2.9)
The precise meaning of (2.9) is given in (2.13). These operators can be extended to bounded linear mappings $\mathcal{W}:H^{1/2,1/4}(\Sigma)\to H^{-1/2,-1/4}(\Sigma)$ and $\mathcal{N}:H^{-1/2,-1/4}(\Sigma)\to H^{1/2,1/4}(\Sigma)$. In addition, $\mathcal{W}$ is coercive, i.e.,
@@
As a matter of fact, [7, Theorem 3.11] even states the following coercivity
@@ 2.3. Weak formulation
Given right-hand sides $f\in L^{2}(Q)$, $u_{0}\in L^{2}(\Omega)$, $u_{D}\in H^{1/2,1/4}(\Sigma)$, and $\phi_{N}\in H^{-1/2,-1/4}(\Sigma)$, we look for weak solutions of (1.1), i.e., $u\in L^{2}(I;H^{1}(\Omega))\cap H^{1}(I;H^{-1}(\Omega))$ and $u^{\rm ext}|_{Q^{\rm ext}_{R}}\in L^{2}(I;H^{1}(\Omega^{\rm ext}_{R}))$ for all $R>0$ such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$. In particular, the PDEs (1.1a)–(1.1b) are to be understood in the distributional sense. Note that $\partial_{t}u^{\rm ext}-\Delta_{\boldsymbol{x}}u^{\rm ext}=0$ readily yields the additional regularity $u^{\rm ext}|_{Q^{\rm ext}_{R}}\in L^{2}(I;H^{1}(\Omega^{\rm ext}_{R}))\cap H^{1}(I;H^{-1}(\Omega^{\rm ext}_{R}))$; see [7, Lemma 2.15]. By the properties of the involved Sobolev spaces mentioned in Section 2.1, the left-hand sides in (1.1) are indeed all well defined.
Let $R>0$ be sufficiently large such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$. With the associated single- and double-layer potential $\widetilde{\mathcal{V}}_{R}$ and $\widetilde{\mathcal{K}}_{R}$ on $\Sigma_{R}$, the representation formula (2.4) applied to $Q^{\rm ext}_{R}$ instead of $Q$ yields that any (non-unique) weak solution of the exterior problem (1.1b)&(1.1d) is given by
@@
In what follows, we show that the radiation condition (1.1g) implies that the last two terms in (2.14) vanish as $R\to\infty$. While this seems to be a fundamental result, we were unable to find the precise statement in the literature and thus give a detailed proof for the sake of completeness. We start with the following auxiliary result, which provides a crude estimate for the gradient $\nabla_{\boldsymbol{x}}u^{\rm ext}$.
@@
By parabolic regularity, e.g., [19, Theorem 1.58], the homogeneous heat equation (1.1b) implies that $u^{\rm ext}$ is even a smooth strong solution of $\partial_{t}u^{\rm ext}-\Delta_{\boldsymbol{x}}u^{\rm ext}=0$ on $Q^{\rm ext}$. In particular, each component of $\nabla_{\boldsymbol{x}}u^{\rm ext}$ satisfies the homogeneous heat equation strongly on $Q^{\rm ext}$. We may thus apply the mean-value property for the heat equation, e.g., [17, Section 2.3, Theorem 3] for each component: Let $R>0$ be sufficiently large such that $\overline{\Omega}\subset B_{R}(\boldsymbol{0})$ and $|u^{\rm ext}(t,{\boldsymbol{x}})|\leq 1$ for all $t\in I$ and all ${\boldsymbol{x}}\in\Omega^{\rm ext}$ with $|{\boldsymbol{x}}|\geq R$. Now, we fix coordinates $t\in I$ and ${\boldsymbol{x}}\in\Omega^{\rm ext}$ with $|{\boldsymbol{x}}|\geq 2R$. For $r>0$, we define the corresponding heat ball
@@
For sufficiently small $r>0$ with $E(t,{\boldsymbol{x}};r)\subset Q^{\rm ext}$, the following mean-value property holds:
@@
If $d=3$, the last integral is just $\tfrac{1}{2\pi}t$, which concludes the proof in this case even for $\epsilon=0$. If $d=2$, we assume without loss of generality that $20<\epsilon<1/$ 2. We write $\ln\frac{t}{4\pi s}=\ln t-\ln 4\pi-\ln s\leq\ln T-\ln s\leq 2\max\{|\ln T|,|\ln s|\}$ and get that
$|\nabla_{\boldsymbol{x}}u^{\rm ext}(t,{\boldsymbol{x}})|$ $\lesssim\frac{1}{t}\int_{0}^{\frac{1}{4\pi}t}s^{-1/2}\Big{(}|\ln T|^{3/2}+|\ln s|^{3/2}\Big{)}\,{\rm d}s$
$=\frac{1}{t}\int_{0}^{\frac{1}{4\pi}t}s^{-1/2-\epsilon}\,s^{\epsilon}\Big{(}|\ln T|^{3/2}+|\ln s|^{3/2}\Big{)}\,{\rm d}s$
$\leq\frac{\max_{s\in[0,T]}s^{\epsilon}\big{(}|\ln T|^{3/2}+|\ln s|^{3/2}\big{)}}{t}\int_{0}^{\frac{1}{4\pi}t}s^{-1/2-\epsilon}\,{\rm d}s\lesssim t^{-1/2-\epsilon}.$
@@
It remains to show that the last two terms in (2.14) tend to zero. To this end, let $(t,{\boldsymbol{x}})\in Q_{R}^{\rm ext}$ be arbitrary but fixed. The simple estimate $|G(s,{\boldsymbol{y}})|\lesssim s^{1/2}|{\boldsymbol{y}}|^{-d-1}$ for all $s\geq 0$, ${\boldsymbol{y}}\in\mathbb{R}^{d}$ (see, e.g., the proof of [7, Theorem 3.11]), the preceding lemma, and $|\Gamma_{R}|\simeq R^{d-1}$ show for sufficiently large $R>0$ and $20<\epsilon<1/$ 2 that
@@
Conversely, in the proof of [7, Theorem 3.11], it is shown that any function of the form
@@
To be more precise, the cited [7, Theorem 3.11] states for $u^{\rm ext}$ defined as in (2.20) that $u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-d})$ and $\nabla_{\boldsymbol{x}}u^{\rm ext}(\cdot,{\boldsymbol{x}})\cdot{\boldsymbol{x}}={\mathcal{O}}(|x|^{-d})$ as $|x|\to\infty$. Using the same argument, one even sees that $u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-\mu})$ and $\nabla_{\boldsymbol{x}}u^{\rm ext}(\cdot,{\boldsymbol{x}})={\mathcal{O}}(|x|^{-\mu})$ as $|x|\to\infty$ for all fixed $\mu\geq 0$ (with hidden constants depending only on $d$, $\mu$, $u^{\rm ext}$, and $T$). The assumption (1.1g) that $u^{\rm ext}$ converges to zero as $|{\boldsymbol{x}}|\to\infty$ thus automatically yields a much stronger decay of $u^{\rm ext}$ and its spatial gradient $\nabla_{\boldsymbol{x}}u^{\rm ext}$. On the other hand, to guarantee that the last two terms in (2.14) converge to zero, it is clear from our proof that one could even allow for a certain growth of $u^{\rm ext}$ (and $\nabla_{\boldsymbol{x}}u^{\rm ext}$). Here and in the following, the minimal required radiation condition would be to explicitly assume that those two terms vanish as $R\to\infty$.